OpenStudy (anonymous):

Find a power series representation for the antiderivative of the integral below integral from 0 to pi of 8sin(x) / x dx

OpenStudy (anonymous):

\[\int_0^\pi8{\sin x\over x}{\rm d}x\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sin(x)=x-\frac{x^3}{3!}+\frac{x^5}{5!}-...\] \[\frac{\sin(x)}{x}=1-\frac{x^2}{3!}+\frac{x^4}{5!}\] at least to three terms

OpenStudy (anonymous):

integrate term by term, then compute

OpenStudy (anonymous):

so I can't substitute the summation equivalent for f(x) = sin(x) ?

OpenStudy (anonymous):

Integrating term by term I got \[\sum_{n=0}^{\Pi} \frac{ (8)(-1)^{2n} (x)^{2n+1} }{ (2n+1)(2n+1)! }\]

OpenStudy (anonymous):

but it's incorrect

OpenStudy (goformit100):

It's not inccorect ... I swear

OpenStudy (anonymous):

okay you have a number to compute, not an anti derivative i assume you are going to compute this as a number, so once you have the power series for \(\frac{\sin(x)}{x}\) and integrate term by term, you have to pick some point at which to stop to evaluate

OpenStudy (anonymous):

the power series you get is not the sum from \(0\) to \(\pi\), it is from \(0\) to \(\infty\)

OpenStudy (anonymous):

i think you have confused two things, one is the power series, one is how to evaluate the integral

OpenStudy (anonymous):

i stopped at 3 terms \[\frac{\sin(x)}{x}=1-\frac{x^2}{3!}+\frac{x^4}{5!}\] then \[\int \frac{\sin(x)}{x}=1-\frac{x^3}{18}+\frac{x^5}{600}\] evaluate at \(\pi\) and \(0\) and subtract

OpenStudy (anonymous):

damn that was wrong \[\int \frac{\sin(x)}{x}dx=x-\frac{x^3}{18}+\frac{x^5}{600}\]

OpenStudy (anonymous):

i see! i didn't integrate from 0 to pi

OpenStudy (anonymous):

yeah you have to compute some number i.e. you are going to make an approximation

OpenStudy (anonymous):

so would the summation be \[\sum_{n=0}^{\infty} \frac{ \pi ^{2n+1} }{(2n+1)(2n+1)! }\]

OpenStudy (anonymous):

i think you are missing a \((-1)^n\) term

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} \frac{(-1)^{n+1} \pi ^{2n+1} }{(2n+1)(2n+1)! }\] might work

OpenStudy (anonymous):

no that wont work

OpenStudy (anonymous):

maybe just \[\sum_{n=0}^{\infty} \frac{(-1)^{n} \pi ^{2n+1} }{(2n+1)(2n+1)! }\]

OpenStudy (anonymous):

yeah, that is the one

OpenStudy (anonymous):

yeah you're right about the (-1)^n. i'll try it and let you know.

OpenStudy (anonymous):

all of that times 8 right?

OpenStudy (anonymous):

i checked, it is right take a look http://www.wolframalpha.com/input/?i= \sum_{n%3D0}^{\infty}+\frac{%28-1%29^{n}+\pi+^{2n%2B1}+}{%282n%2B1%29%282n%2B1%29!+} compare to http://www.wolframalpha.com/input/?i=intgral+0+to+pi+sin%28x%29%2Fx

OpenStudy (anonymous):

yes, times 8

OpenStudy (anonymous):

yes! it's correct!

OpenStudy (anonymous):

thank you so much!